# Rational solutions of Painlevé V from Hankel determinants and the asymptotics of their pole locations

### v2

@@ 5.2.1 Second transformation, model problem and its solution: the case $\mathrm{Re}\,(s)>0$
$\lim_{z\to\infty}\begin{bmatrix}\boldsymbol{\phi}\\ \boldsymbol{\psi}\end{bmatrix}$ $=\begin{bmatrix}\frac{\vartheta\left(G-\frac{\tau+1}{2}\right)}{\omega_{1}\vartheta^{\prime}\left(-\frac{\tau+1}{2}\right)}{\rm e}^{i\pi K\mathfrak{u}(\infty)}&0\\ 0&\frac{\vartheta\left(G-\frac{\tau+1}{2}\right)}{\omega_{1}\vartheta^{\prime}\left(-\frac{\tau+1}{2}\right)}{\rm e}^{-i\pi K\mathfrak{u}(\infty)}\end{bmatrix}.$ (5.97)
@@
$\widehat{M}(z;G,K,\infty):=\frac{\omega_{1}\vartheta^{\prime}\left(-\frac{\tau+1}{2}\right){\rm e}^{-i\pi K\mathfrak{u}(\infty)\sigma_{3}}}{\vartheta\left(G-\frac{\tau+1}{2}\right)}\begin{bmatrix}\boldsymbol{\phi}(z;\infty)\\ \boldsymbol{\psi}(z;\infty)\end{bmatrix}.$ (5.98)
Remark 5.2
We have emphasized in the notation that $\infty$ as a point plays a role in the expression. We will use later the same formula, but replacing $\infty$ with the point $z=0$. It is important to point out that, irrespectively of what point we replace instead of $\infty$, the matrix satisfies the same boundary relations (5.100) below, and also the same behaviour near the branch-points $z=1,b,a_{\pm}$. The only difference is that if we replace $\infty$ by a point $z_{0}$, the matrix $\widehat{M}(z;G,K,z_{0})$ will then vanish at infinity and have a simple pole at $z=z_{0}$ with singular part proportional to the identity matrix:
$\lim_{z\to\infty}\widehat{M}(z;G,K,z_{0})={\boldsymbol{0}},\qquad\widehat{M}(z;G,K,z_{0})=\frac{\omega_{1}h(0)}{\mathfrak{u}^{\prime}(z_{0})}\frac{\mathbf{1}}{z-z_{0}}+\mathcal{O}(1),\ \ z\to z_{0}.$ (5.99)
We are going to use this observation later. $\triangle$
It is a direct verification using the definition of the Abel map (5.88), the properties of the theta function (5.94), and the jump relation of $h$ (5.85), that each row $\boldsymbol{\phi},\boldsymbol{\psi}$ satisfies the three boundary value relations similar to those in (5.81):
$\widehat{M}(z_{+})$ $=\widehat{M}(z_{-})\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ $z\in\Gamma_{a_{-}}^{b}$
$\widehat{M}(z_{+})$ $=\widehat{M}(z_{-})\begin{bmatrix}0&-{\rm e}^{i\pi K}\\ {\rm e}^{-i\pi K}&0\end{bmatrix}$ $z\in\Gamma_{b}^{a_{+}}$
$\widehat{M}(z_{+})$ $=\widehat{M}(z_{-})\begin{bmatrix}0&-{\rm e}^{-2i\pi\left(G-\frac{K\tau}{2}\right)}\\ {\rm e}^{2i\pi\left(G-\frac{K\tau}{2}\right)}&0\end{bmatrix}$ $z\in\Gamma_{b}^{1}.$ (5.100)
By matching these to the boundary relations (5.81) we deduce that
Proposition 5.3
The solution of the RHP (5.9) is given by $M(z)=\widehat{M}(z;G,K,\infty)$ in (5.98) with the values of the constants $G,K$ given by
$G=\frac{1}{2i\pi}\left(-n\Omega_{1}-\nu+n\Omega_{2}\tau\right)+\frac{\tau+1}{2},\qquad K=\frac{n\Omega_{2}}{i\pi}+1.$ (5.101)
Proof. Comparing the boundary relations (5.100) and (5.81) we obtain the system
$\left\{\begin{array}[]{l}\displaystyle G-\frac{K\tau}{2}=-\frac{n\Omega_{1}+\nu}{2i\pi}+\frac{1}{2}\\[10.0pt] \displaystyle K=\frac{n\Omega_{2}}{i\pi}+1,\end{array}\right.$ (5.102)
from which the relations follow. $\blacksquare$
The solvability of the RHP 5.9 and hence of the RHP 5.8 depends entirely on the non-vanishing of the expression
$\vartheta\left(\frac{n\Omega_{1}+\nu-n\Omega_{2}\tau}{2i\pi}\right)$ (5.103)
@@
$(H_{1},H_{2})\in{\mathbb{Z}}^{2}$
$H_{1}:=\frac{n\Omega_{1}}{2i\pi}+\mathrm{Re}\,\left(\frac{\nu}{2i\pi}\right)-\frac{\mathrm{Re}\,\tau}{\mathrm{Im}\,\tau}\mathrm{Im}\,\left(\frac{\nu}{2i\pi}\right)+\frac{1}{2}$
$H_{2}:=\frac{n\Omega_{2}}{2i\pi}-\frac{\mathrm{Im}\,\left(\frac{\nu}{2i\pi}\right)}{\mathrm{Im}\,\tau}+\frac{1}{2}.$ (5.104)
Observe that both $H_{1},H_{2}$ are real functions of $s$ since $\Omega_{1},\Omega_{2}\in i{\mathbb{R}}$. The modular parameter $\tau$ is also, in a very implicit way, a function of $s$ since the branch-points of the radical $R$ (5.60) are determined by the Boutroux conditions (4.31). Ditto for $\nu$, which depends on $s$ implicitly via the formula (5.68).
Remark 5.3 (Explanation of Fig. 3.)
The two quantization conditions (5.2.1) can be interpreted as describing a mesh of level sets of the two functions $H_{1},H_{2}$ (both functions of the complex parameter $s$); in the various panels in Fig. 3 these are precisely forming the mesh of curves that populate the interior of $EoT$. The intersection points of this mesh are the points where the quantization conditions (5.2.1) hold, and hence where the solution of the model problem RHP 5.8 ceases to exist. They also approximate very precisely the zeros of the rational solutions (in fact better than expected), with only obvious deviations near the boundaries of the two halves of $EoT$, since there the elliptic curve of the radical $R$ (5.60) degenerates (i.e., two branch-points come together). $\triangle$
5.2.2 Second transformation, model problem and its solution: the case $\mathrm{Re}\,(s)<0$
With reference to Fig. 8, right pane, we denote by $\Lambda_{\pm}^{(x,y)}$ the region bounded by $\Gamma_{x}^{y}$ and ±$\mathscr{L}_{\pm}^{(x,y)}$, for $x,y\in\{a_{+},a_{-},b,1\}$ and we call them lens regions (as opposed to their boundaries, which we refer to as the lens arcs).
Definition of $T_{0}$ for $\mathrm{Re}\,(s)<0$.
We refer to the right pane of Fig. 8 and then set
$T_{0}(z):=\left\{\begin{array}[]{lc}\displaystyle W(z)\begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix}&z\in\Lambda_{-}^{(a_{+},a_{-})}\\[10.0pt] W(z)\begin{bmatrix}1&0\\ -\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix}&z\in\Lambda_{+}^{(a_{+},a_{-})}\end{array}\right.$ (5.105)
while, for the lens $\Lambda_{\pm}^{(b,1)}$ adjacent to $\Gamma_{b}^{1}$ we have the same as in (5.45). The construction is similar to the previous case, but with the notable difference that now the main arcs $\Gamma_{m}$ consists of two disjoint arcs (see Fig. 8, right pane)
$\Gamma_{m}=\Gamma_{b}^{1}\cup\Gamma_{a_{+}}^{a_{-}},\qquad\Gamma_{c}=\Gamma_{b}^{a_{+}}\cup\Gamma_{a_{-}}^{b}.$ (5.106)
@@ 5.2.2 Second transformation, model problem and its solution: the case $\mathrm{Re}\,(s)<0$
$\varphi(z;s):=\int_{a_{-}}^{z}\frac{2}{w^{2}}\sqrt{w^{2}+\frac{s^{2}}{4}+\frac{Aw^{2}}{w-1}}{\mathrm{d}}w,$ (5.107)
where now, however, the branch cuts of the radical in the integrand consist of the two arcs $\Gamma_{b}^{1}$, $\Gamma_{a_{+}}^{a_{-}}$ of Fig. 8, right pane, and $\varphi$ has additionally a branch cut $[1,\infty)$ and another one running along $\Gamma_{b}^{a_{+}}$. Now the effective potential satisfies the following boundary value relations:
$\varphi(z_{+})+\varphi(z_{-})$ $=0$ $z\in\Gamma_{a_{+}}^{a_{-}}$
$\varphi(z_{+})+\varphi(z_{-})$ $=2\Omega_{1}$ $z\in\Gamma_{b}^{1}$
$\varphi(z_{+})-\varphi(z_{-})$ $=0$ $z\in\Gamma_{a_{-}}^{b}$
$\varphi(z_{+})-\varphi(z_{-})$ $=2\Omega_{2}$ $z\in\Gamma^{a_{+}}_{b}$
$\varphi(z_{+})-\varphi(z_{-})$ $=4i\pi$ $z\in[1,\infty)$ (5.108)
@@
$\Omega_{1}:=\int_{a_{-}}^{b}\varphi^{\prime}(w){\mathrm{d}}w,\ \ \ \ \Omega_{2}=\int_{a_{-}}^{a_{+}}\varphi^{\prime}(w_{+}){\mathrm{d}}w.$ (5.109)
A direct computation shows that the matrix $T_{0}(z)$ satisfies the following RHP
Problem 5.10
@@
$T_{0}(z_{+})=T_{0}(z_{-})J_{T_{0}}(z),\ \ \ \ z\in\gamma$ (5.110)
$T_{0}(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{z}\right),\ \ \ \ z\to\infty$ (5.111)
where (we set for brevity $Q_{\pm}:=Q(z_{\pm})$, $\varphi_{\pm}:=\varphi(z_{\pm})$ below)
$J_{T_{0}}(z)=\left\{\begin{array}[]{lc}J_{{}_{W}}(z)=\begin{bmatrix}{\rm e}^{-n\Omega_{2}}&Q{\rm e}^{\frac{n}{2}(\varphi_{+}+\varphi_{-})}\\ 0&{\rm e}^{n\Omega_{2}}\end{bmatrix}&z\in\Gamma_{b}^{a_{+}}\\[12.0pt] J_{{}_{W}}(z)=\begin{bmatrix}1&Q{\rm e}^{n\varphi}\\ 0&1\end{bmatrix}&z\in\Gamma^{b}_{a_{-}}\\[12.0pt] \begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi}}{Q}&1\end{bmatrix}&z\in\mathscr{L}_{+}^{(a_{+},a_{-})}\cup\mathscr{L}_{-}^{(a_{+},a_{-})}\\[15.0pt] \begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi}}{\varkappa Q}&1\end{bmatrix}&z\in\mathscr{L}_{+}^{(b,1)}\\[15.0pt] \begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi-2i\pi{{\rho}}}}{\varkappa Q}&1\end{bmatrix}&z\in\mathscr{L}_{-}^{(b,1)}\\[15.0pt] \begin{bmatrix}0&\varkappa Q_{+}{\rm e}^{n\Omega_{1}}\\ -\frac{{\rm e}^{-n\Omega_{1}}}{\varkappa Q_{+}}&0\end{bmatrix}&z\in\Gamma_{b}^{1}\\[15.0pt] \begin{bmatrix}0&Q\\ -\frac{1}{Q}&0\end{bmatrix}&z\in\Gamma_{a_{+}}^{a_{-}}.\end{array}\right.$ (5.112)
Problem 5.11
@@
$M_{{}_{Q}}(z_{+})$ $=M_{{}_{Q}}(z_{-})$ $z\in\Gamma_{a_{-}}^{b}\cup[1,\infty)$
$M_{{}_{Q}}(z_{+})$ $=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&Q(z)\\ \frac{-1}{Q(z)}&0\end{bmatrix}$ $z\in\Gamma_{a_{+}}^{a_{-}}$
$M_{{}_{Q}}(z_{+})$ $=M_{{}_{Q}}(z_{-})\begin{bmatrix}{\rm e}^{-n\Omega_{2}}&0\\ 0&{\rm e}^{n\Omega_{2}}\end{bmatrix}$ $z\in\Gamma_{b}^{a_{+}}$
$M_{{}_{Q}}(z_{+})$ $=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&\varkappa Q(z_{+}){\rm e}^{n\Omega_{1}}\\ \frac{-{\rm e}^{-n\Omega_{1}}}{\varkappa Q(z_{+})}&0\end{bmatrix}$ $z\in\Gamma_{b}^{1}$ (5.113)
with $\varkappa$ as in (5.39) and $Q$ as in (5.3). Furthermore the following local behaviours hold
$M_{{}_{Q}}(z)$ $=\mathbf{1}+\mathcal{O}(z^{-1})$ $|z|\to\infty$ (5.114)
$M_{{}_{Q}}(z)$ $=\mathcal{O}\left(\frac{1}{(z-q)^{\frac{1}{4}}}\right),\ \ \ q=1,b,a_{\pm}$ $z\to q.$ (5.115)
Like before, to further normalize the problem we need to construct an appropriate Szegö function.
The Szegö function.
Let $R(z)$ denote the radical function (5.60), but now with the branch cuts of the radical chosen along $\Gamma_{m}=\Gamma_{b}^{1}\cup\Gamma_{a_{+}}^{a_{-}}$ (right pane of Fig. 8). Consider the following expression:
$S(z)=R(z)\left(\int_{b}^{1}\frac{\ln Q(w_{+})+\ln\varkappa-\nu}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2i\pi}+\int_{a_{+}}^{a_{-}}\frac{\ln Q(w)}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2i\pi}\right).$ (5.116)
Following analogous considerations as in the case $\mathrm{Re}\,(s)>0$ (Sec. 5.2.1) we finally obtain the same expression (5.67), where the condition that $S(z)$ is bounded at infinity imposes the same constraint on $\nu$ as in (5.68). Similarly to Prop. 5.2 we have now:
Proposition 5.4 (Szegö function for $\mathrm{Re}\,(s)<0$)
@@
$S(z_{+})+S(z_{-})=\left\{\begin{array}[]{lc}\ln\left(\varkappa Q(z_{+})\right)-\nu&z\in\Gamma_{b}^{1}\\ \ln Q(z)&z\in\Gamma_{a_{+}}^{a_{-}}.\end{array}\right.$ (5.119)
Furthermore $S(z)$ is bounded near $z=b,a_{\pm}$ and near $z=1$ it has the behaviour
$S(z)=-\frac{{\rho}}{2}\ln(z-1)+\mathcal{O}(1).$ (5.120)
With this new Szegö function we normalize RHP 5.11 and define the new model problem
$M(z):={\rm e}^{S(\infty)\sigma_{3}}M_{{}_{Q}}(z){\rm e}^{-S(z)\sigma_{3}}.$ (5.121)
@@
Problem 5.12
Find a matrix–valued function $M(z)$ analytic in ${\mathbb{C}}\setminus\Gamma_{m}$ such that
$M(z_{+})$ $=M(z_{-})\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ $z\in\Gamma_{a_{+}}^{a_{-}}$
$M(z_{+})$ $=M(z_{-})\begin{bmatrix}{\rm e}^{-n\Omega_{2}}\\ 0&{\rm e}^{n\Omega_{2}}\end{bmatrix}$ $z\in\Gamma_{b}^{a_{+}}$
$M(z_{+})$ $=M(z_{-})\begin{bmatrix}0&{\rm e}^{n\Omega_{1}+\nu}\\ -{\rm e}^{-n\Omega_{1}-\nu}&0\end{bmatrix}$ $z\in\Gamma_{b}^{1}$ (5.122)

### v3

@@ 5.2.1 Second transformation, model problem and its solution: The case $\boldsymbol{\operatorname{Re}(s)>0}$
$\lim_{z\to\infty}\begin{bmatrix}\boldsymbol{\phi}\\ \boldsymbol{\psi}\end{bmatrix}=\begin{bmatrix}\displaystyle\frac{\vartheta\left(G-\frac{\tau+1}{2}\right)}{\omega_{1}\vartheta^{\prime}\left(-\frac{\tau+1}{2}\right)}{\rm e}^{{\rm i}\pi K\mathfrak{u}(\infty)}&0\\ 0&\displaystyle\frac{\vartheta\left(G-\frac{\tau+1}{2}\right)}{\omega_{1}\vartheta^{\prime}\left(-\frac{\tau+1}{2}\right)}{\rm e}^{-{\rm i}\pi K\mathfrak{u}(\infty)}\end{bmatrix}$.
@@
$\widehat{M}(z;G,K,\infty):=\frac{\omega_{1}\vartheta^{\prime}\left(-\frac{\tau+1}{2}\right){\rm e}^{-{\rm i}\pi K\mathfrak{u}(\infty)\sigma_{3}}}{\vartheta\left(G-\frac{\tau+1}{2}\right)}\begin{bmatrix}\boldsymbol{\phi}(z;\infty)\\ \boldsymbol{\psi}(z;\infty)\end{bmatrix}$. (5.37)
Remark 5.17.
We have emphasized in the notation that $\infty$ as a point plays a role in the expression. We will use later the same formula, but replacing $\infty$ with the point $z=0$. It is important to point out that, irrespectively of what point we replace instead of $\infty$, the matrix satisfies the same boundary relations (5.38) below, and also the same behaviour near the branch-points $z=1,b,a_{\pm}$}. The only difference is that if we replace $\infty$ by a point $z_{0}$, the matrix $\smash{\widehat{M}(z;G,K,z_{0})}$ will then vanish at infinity and have a simple pole at $z=z_{0}$ with singular part proportional to the identity matrix
$\lim_{z\to\infty}\widehat{M}(z;G,K,z_{0})={\boldsymbol{0}},\qquad\widehat{M}(z;G,K,z_{0})=\frac{\omega_{1}h(0)}{\mathfrak{u}^{\prime}(z_{0})}\frac{\mathbf{1}}{z-z_{0}}+\mathcal{O}(1),\qquad z\to z_{0}.$
We are going to use this observation later.
It is a direct verification using the definition of the Abel map (5.33), the properties of the theta function (5.34), and the jump relation of $h$ (5.31), that each row $\boldsymbol{\phi}$, $\boldsymbol{\psi}$ satisfies the three boundary value relations similar to those in (5.27)
$\widehat{M}(z_{+})=\widehat{M}(z_{-})\begin{bmatrix}0&1\\ -1&0\end{bmatrix},\qquad z\in\Gamma_{a_{-}}^{b},$
$\widehat{M}(z_{+})=\widehat{M}(z_{-})\begin{bmatrix}0&-{\rm e}^{{\rm i}\pi K}\\ {\rm e}^{-{\rm i}\pi K}&0\end{bmatrix},\qquad z\in\Gamma_{b}^{a_{+}},$
$\widehat{M}(z_{+})=\widehat{M}(z_{-})\begin{bmatrix}0&-{\rm e}^{-2{\rm i}\pi(G-\frac{K\tau}{2})}\\ {\rm e}^{2{\rm i}\pi(G-\frac{K\tau}{2})}&0\end{bmatrix},\qquad z\in\Gamma_{b}^{1}.$ (5.38)
By matching these to the boundary relations (5.27), we deduce the following.
Proposition 5.18.
The solution of the Riemann–Hilbert Problem 5.15 is given by $M(z)=\widehat{M}(z;G,K,\infty)$ in (5.37) with the values of the constants $G$, $K$ given by
$G=\frac{1}{2{\rm i}\pi}\left(-n\Omega_{1}-\nu+n\Omega_{2}\tau\right)+\frac{\tau+1}{2},\qquad K=\frac{n\Omega_{2}}{{\rm i}\pi}+1.$ (5.39)
Proof.
Comparing the boundary relations (5.38) and (5.27), we obtain the system
$G-\frac{K\tau}{2}=-\frac{n\Omega_{1}+\nu}{2{\rm i}\pi}+\frac{1}{2},\qquad\displaystyle K=\frac{n\Omega_{2}}{{\rm i}\pi}+1,$
from which the relations follow. ∎
The solvability of the Riemann–Hilbert Problem 5.15 and hence of the Riemann–Hilbert Problem 5.13 depends entirely on the non-vanishing of the expression
$\vartheta\left(\frac{n\Omega_{1}+\nu-n\Omega_{2}\tau}{2{\rm i}\pi}\right)$ (5.40)
@@
$(H_{1},H_{2})\in{\mathbb{Z}}^{2},$
$H_{1}:=\frac{n\Omega_{1}}{2{\rm i}\pi}+\operatorname{Re}\left(\frac{\nu}{2{\rm i}\pi}\right)-\frac{\operatorname{Re}\tau}{\operatorname{Im}\tau}\operatorname{Im}\left(\frac{\nu}{2{\rm i}\pi}\right)+\frac{1}{2},$
$H_{2}:=\frac{n\Omega_{2}}{2{\rm i}\pi}-\frac{\operatorname{Im}\left(\frac{\nu}{2{\rm i}\pi}\right)}{\operatorname{Im}\tau}+\frac{1}{2}.$ (5.41)
Observe that both $H_{1}$, $H_{2}$ are real functions of $s$ since $\Omega_{1},\Omega_{2}\in{\rm i}{\mathbb{R}}$. The modular parameter $\tau$ is also, in a very implicit way, a function of $s$ since the branch-points of the radical $R$ (5.19) are determined by the Boutroux conditions (4.9). Ditto for $\nu$, which depends on $s$ implicitly via the formula (5.22).
Remark 5.19 (explanation of Figure 3.).
The two quantization conditions (5.41) can be interpreted as describing a mesh of level sets of the two functions $H_{1}$, $H_{2}$ (both functions of the complex parameter $s$). In the various panels in Figure 3 these are precisely forming the mesh of curves that populate the interior of $\rm EoT$}. The intersection points of this mesh are the points where the quantization conditions (5.41) hold, and hence where the solution of the model problem Riemann–Hilbert Problem 5.13 ceases to exist. They also approximate very precisely the zeros of the rational solutions (in fact better than expected), with only obvious deviations near the boundaries of the two halves of $\rm EoT$}, since there the elliptic curve of the radical $R$ (5.19) degenerates (i.e., two branch-points come together).
5.2.2 Second transformation, model problem and its solution: The case $\boldsymbol{\operatorname{Re}(s)<0}$
With reference to Figure 8, right pane, we denote by $\Lambda_{\pm}^{(x,y)}$ the region bounded by $\Gamma_{x}^{y}$ and $\mathscr{L}_{\pm}^{(x,y)}$, for $x,y\in\{a_{+},a_{-},b,1\}$ and we call them lens regions (as opposed to their boundaries, which we refer to as the lens arcs).
Definition of $\boldsymbol{T_{0}}$ for $\boldsymbol{\operatorname{Re}(s)<0}$. We refer to the right pane of Figure 8 and then set
$T_{0}(z):=\begin{cases}\displaystyle W(z)\begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix},&z\in\Lambda_{-}^{(a_{+},a_{-})},\vskip 2.84526pt\\ W(z)\begin{bmatrix}1&0\\ \displaystyle-\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix},&z\in\Lambda_{+}^{(a_{+},a_{-})},\end{cases}$
while, for the lens $\Lambda_{\pm}^{(b,1)}$ adjacent to $\Gamma_{b}^{1}$ we have the same as in (5.11). The construction is similar to the previous case, but with the notable difference that now the main arcs $\Gamma_{m}$ consists of two disjoint arcs (see Figure 8, right pane)
$\Gamma_{m}=\Gamma_{b}^{1}\cup\Gamma_{a_{+}}^{a_{-}},\qquad\Gamma_{c}=\Gamma_{b}^{a_{+}}\cup\Gamma_{a_{-}}^{b}.$
@@ 5.2.2 Second transformation, model problem and its solution: The case $\boldsymbol{\operatorname{Re}(s)<0}$
$\varphi(z;s):=\int_{a_{-}}^{z}\frac{2}{w^{2}}\sqrt{w^{2}+\frac{s^{2}}{4}+\frac{Aw^{2}}{w-1}}{\mathrm{d}}w,$
where now, however, the branch cuts of the radical in the integrand consist of the two arcs $\Gamma_{b}^{1}$, $\Gamma_{a_{+}}^{a_{-}}$ of Figure 8, right pane, and $\varphi$ has additionally a branch cut $1,\infty)$ and another one running along $\Gamma_{b}^{a_{+}}$. Now the effective potential satisfies the following boundary value relations:
$\varphi(z_{+})+\varphi(z_{-})=0,\qquad z\in\Gamma_{a_{+}}^{a_{-}},$
$\varphi(z_{+})+\varphi(z_{-})=2\Omega_{1},\qquad z\in\Gamma_{b}^{1},$
$\varphi(z_{+})-\varphi(z_{-})=0,\qquad z\in\Gamma_{a_{-}}^{b},$
$\varphi(z_{+})-\varphi(z_{-})=2\Omega_{2},\qquad z\in\Gamma^{a_{+}}_{b}$,
$\varphi(z_{+})-\varphi(z_{-})=4{\rm i}\pi,\qquad z\in[1,\infty),$
@@
$\Omega_{1}:=\int_{a_{-}}^{b}\varphi^{\prime}(w){\mathrm{d}}w,\qquad\Omega_{2}=\int_{a_{-}}^{a_{+}}\varphi^{\prime}(w_{+}){\mathrm{d}}w.$ (5.42)
A direct computation shows that the matrix $T_{0}(z)$ satisfies the following Riemann–Hilbert problem.
Problem 5.20.
@@
$T_{0}(z_{+})=T_{0}(z_{-})J_{T_{0}}(z),\qquad z\in\gamma,\qquad T_{0}(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{z}\right),\qquad z\to\infty,$
where ((we set for brevity $Q_{\pm}:=Q(z_{\pm})$, $\varphi_{\pm}:=\varphi(z_{\pm})$ below))
$J_{T_{0}}(z)=\begin{cases}J_{{}_{W}}(z)=\begin{bmatrix}{\rm e}^{-n\Omega_{2}}&Q{\rm e}^{\frac{n}{2}(\varphi_{+}+\varphi_{-})}\\ 0&{\rm e}^{n\Omega_{2}}\end{bmatrix},&z\in\Gamma_{b}^{a_{+}},\\[10.0pt] J_{{}_{W}}(z)=\begin{bmatrix}1&Q{\rm e}^{n\varphi}\\ 0&1\end{bmatrix},&z\in\Gamma^{b}_{a_{-}},\\[10.0pt] \begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi}}{Q}&1\end{bmatrix},&z\in\mathscr{L}_{+}^{(a_{+},a_{-})}\cup\mathscr{L}_{-}^{(a_{+},a_{-})},\\[10.0pt] \begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi}}{\varkappa Q}&1\end{bmatrix},&z\in\mathscr{L}_{+}^{(b,1)},\\[10.0pt] \begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi-2{\rm i}\pi{{\rho}}}}{\varkappa Q}&1\end{bmatrix},&z\in\mathscr{L}_{-}^{(b,1)},\\[10.0pt] \begin{bmatrix}0&\varkappa Q_{+}{\rm e}^{n\Omega_{1}}\\ \displaystyle-\frac{{\rm e}^{-n\Omega_{1}}}{\varkappa Q_{+}}&0\end{bmatrix},&z\in\Gamma_{b}^{1},\\[10.0pt] \begin{bmatrix}0&Q\\ -\frac{1}{Q}&0\end{bmatrix},&z\in\Gamma_{a_{+}}^{a_{-}}.\end{cases}$
Problem 5.21.
@@
$M_{{}_{Q}}(z_{+})=M_{{}_{Q}}(z_{-}),\qquad z\in\Gamma_{a_{-}}^{b}\cup[1,\infty),$
$M_{{}_{Q}}(z_{+})=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&Q(z)\\ \displaystyle\frac{-1}{Q(z)}&0\end{bmatrix},\qquad z\in\Gamma_{a_{+}}^{a_{-}},$
$M_{{}_{Q}}(z_{+})=M_{{}_{Q}}(z_{-})\begin{bmatrix}{\rm e}^{-n\Omega_{2}}&0\\ 0&{\rm e}^{n\Omega_{2}}\end{bmatrix},\qquad z\in\Gamma_{b}^{a_{+}},$
$M_{{}_{Q}}(z_{+})=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&\varkappa Q(z_{+}){\rm e}^{n\Omega_{1}}\\ \displaystyle\frac{-{\rm e}^{-n\Omega_{1}}}{\varkappa Q(z_{+})}&0\end{bmatrix},\qquad z\in\Gamma_{b}^{1},$
with $\varkappa$ as in (5.10) and $Q$ as in (5.1). Furthermore, the following local behaviours hold:
$M_{{}_{Q}}(z)=\mathbf{1}+\mathcal{O}\bigl(z^{-1}\bigr),\qquad|z|\to\infty,$
$M_{{}_{Q}}(z)=\mathcal{O}\left(\frac{1}{(z-q)^{\frac{1}{4}}}\right),\qquad q=1,b,a_{\pm},\qquad z\to q.$
Like before, to further normalize the problem we need to construct an appropriate Szegő function.
The Szegő function. Let $R(z)$ denote the radical function (5.19), but now with the branch cuts of the radical chosen along $\Gamma_{m}=\Gamma_{b}^{1}\cup\Gamma_{a_{+}}^{a_{-}}$ (right pane of Figure 8). Consider the following expression:
$S(z)=R(z)\left(\int_{b}^{1}\frac{\ln Q(w_{+})+\ln\varkappa-\nu}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2{\rm i}\pi}+\int_{a_{+}}^{a_{-}}\frac{\ln Q(w)}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2{\rm i}\pi}\right).$ (5.43)
Following analogous considerations as in the case $\operatorname{Re}(s)>0$ (see Section 5.2.1), we finally obtain the same expression (5.21), where the condition that $S(z)$ is bounded at infinity imposes the same constraint on $\nu$ as in (5.22). Similarly to Proposition 5.14, we have now the following.
Proposition 5.22 (Szegő function for $\operatorname{Re}(s)<0$).
@@
$S(z_{+})+S(z_{-})=\begin{cases}\ln\left(\varkappa Q(z_{+})\right)-\nu,&z\in\Gamma_{b}^{1},\\ \ln Q(z),&z\in\Gamma_{a_{+}}^{a_{-}}.\end{cases}$
Furthermore, $S(z)$ is bounded near $z=b,a_{\pm}$ and near $z=1$ it has the behaviour
$S(z)=-\frac{{\rho}}{2}\ln(z-1)+\mathcal{O}(1).$
With this new Szegő function, we normalize Riemann–Hilbert Problem 5.21 and define the new model problem
$M(z):={\rm e}^{S(\infty)\sigma_{3}}M_{{}_{Q}}(z){\rm e}^{-S(z)\sigma_{3}}.$
@@
Problem 5.23.
Find a matrix-valued function $M(z)$ analytic in ${\mathbb{C}}\setminus\Gamma_{m}$ such that
$M(z_{+})=M(z_{-})\begin{bmatrix}0&1\\ -1&0\end{bmatrix},\qquad z\in\Gamma_{a_{+}}^{a_{-}},$
$M(z_{+})=M(z_{-})\begin{bmatrix}{\rm e}^{-n\Omega_{2}}\\ 0&{\rm e}^{n\Omega_{2}}\end{bmatrix},\qquad z\in\Gamma_{b}^{a_{+}},$
$M(z_{+})=M(z_{-})\begin{bmatrix}0&{\rm e}^{n\Omega_{1}+\nu}\\ -{\rm e}^{-n\Omega_{1}-\nu}&0\end{bmatrix},\qquad z\in\Gamma_{b}^{1}$ (5.44)